# Understanding Retrieval Robustness for Retrieval-Augmented Image Captioning

### v1

@@ 4.4 Attention and Model Behavior
Finally, we visualize the self-attention and cross-attention to locate the heads and layers in the model structure that may contribute to the majority voting mechanism within the decoding stage. This is crucial because all interactions between captions (self-attention) and images (cross-attention) take place in this stage.
@@
To comprehensively investigate the distributional nuances across various input components in attention, we partition the textual prompt into five distinct segments: begin of the sentence token (<BOS>), prompt tokens before retrieved captions (prefix), i.e. “Similar image show”, retrieved captions (retrieval) $cap_{1},\cdots cap_{k}$, prompt tokens before generation (suffix), i.e. “This image show”, and the generation itself. For image patches, we segment them into two parts – the CLS patch and all other patches.
Let $P_{k}$ denote the sets of indices, where $k=1,2,\ldots,5$ for five partitions. For text prompts, each $P_{k}$ corresponds to the indices of the tokens in each segment. To track the occurrence of max attention values in $P_{k}$, we define the indicator function $\mathbbm{1}_{I_{k}(i,j)}$ as follows:
$\mathbbm{1}_{I_{k}(i,j)}=\begin{cases}1&\text{if }\arg\max_{z}Atten(j,z)_{i}\in P_{k}\\ 0&\text{otherwise}\end{cases},$ (2)
where $\arg\max_{z}Atten(j,z)_{i}$ denotes the index of the context token with the maximum attention score. For self-attention, regarding each sample $i$, $Atten(j,z)$ represents the attention score between the $j^{th}$ generated token and the $z^{th}$ text context token, denoted as $Atten(j,z)=SA_{text}(j,z)$.
For cross-attention, regarding each sample $i$, we do both text-centric or an image-centric analysis. The text-centric analysis, denoted as $Atten(j,z)=XA_{text}(j,z)$, measures attention between the $j^{th}$ image patch and the $z^{th}$ text token, helping identify which segments of the text receive the highest cross-attention scores in relation to the image patches. Alternatively, for the two segments of image patches, we can reuse the $P_{k}$ notation, where the index of the image patches are partitioned into $P_{1}$ and $P_{2}$, representing the CLS patch and other patches, respectively. Then the image-centric analysis, denoted as $Atten(j,z)=XA_{img}(j,z)$ is the attention between the $j^{th}$ generated token and the $z^{th}$ image patch in sample $i$, which helps identify if the $P_{1}$ (the CLS patch) receives the highest cross-attention scores in relation to the generated tokens.
Then we can calculate the proportion of occurrences of the maximum score in $P_{k}$ by averaging through all generated tokens for the entire dataset.
@@
We gather attention scores between the generated tokens and context tokens, and categorize the distribution of the maximum self-attention into the five text partitions (Equation 2).
Figure 5(a) illustrates the dynamic changes of the distribution of maximum self-attention scores as the layer depth increases. Notably, at the initial layers, a majority of attention heads exhibit heightened focus on retrieved captions or the current context for generation. However, after the second layer, we observe an increased emphasis on the beginning of sentence token (</s>). This behavior is consistent with prior research on the attention mechanism of GPT-2 Vig and Belinkov (2019). See the similar tendency for SmallCap in Figure LABEL:subfig:gpt2-self-attention.
@@
First, similar to the self-attention, we categorize the occurrence of the maximum cross-attention to the five text partitions. As shown in Figure 5(b), it is obvious that across nearly all attention heads, the cross-attention attains its maximum value between the image and the retrieved captions or between the image and the current generation.
Second, we locate image patches that the text predominantly focuses on, i.e. the CLS (class) patch or the remaining patches. In Figure 5(c), we can observe that not until in the final layers of the model, the generated tokens do not pay strong attention to the non-CLS patches. Overall, these observation suggests that the model attends to both modalities during the caption generation process. However the lacking of strong attention to non-CLS patches might give a clue of why the model gets misled by text prompts, even when irrelevant information is absent in the provided image.
@@ 5 Improving Robustness to Retrieval via Sampling
Figure 6: Qualitative examples of generated captions when randomly retrieving four captions for a given image.
To improve the resilience of the model to noisy tokens obtained from retrieval, instead of training with the top-$k$ retrieved captions, we randomly sample $k$ retrieved captions from a more extensive list for a given image. In this configuration, the model learns from more diverse context that includes both top relevant and lower-ranked captions.
@@ 5.1 Experimental Setup
We sample $k$ captions randomly from the top-$N$=7 retrieved captions during training [^4] We sample from the top-$N$=7 for alignment with the baseline; see the Appendix for an ablation on varying $N$.. Following Ramos et al. (2023b), we train smallCap with the OPT-350m decoder on the COCO captioning dataset (Chen et al., 2015) for 10 epochs on a NVIDIA A100 GPU with the default learning rate of 1e-4 and batch size of 64. We experiment with $k$ ranges from one to four.
COCO Eval
Train $k$ top-$k$ last-$k$ random
top-k 1 115.1 112.2 73.2
sample-k 1 116.0 $+0.9$ 115.0 $+2.8$ 98.9 $+25.7$
top-k 2 116.8 115.0 67.4
sample-k 2 117.4 $+0.6$ 116.8 $+1.8$ 84.6 $+16.8$
top-k 3 118.3 117.1 71.8
sample-k 3 118.5 $+0.2$ 117.3 $+0.2$ 77.6 $+5.8$
top-k 4 120.1 117.1 70.1
sample-k 4 119.2 $-0.9$ 118.6 $+1.5$ 73.1 $+3.0$
c-sample-k 4 119.3 $-0.8$ 118.9 $+1.8$ 72.6 + 2.5
Table 3: Top-$k$ and sample-$k$ evaluated on COCO validation set. We report CIDEr scores. Training with sampling the retrieved captions almost always outperforms SmallCap for all $k$ values. It also reduces the gap between using top-relevant and low-ranked retrieved captions. Results are averaged over three seeds. Improved scores are in bold.
@@ 5.2 Evaluation and Results
In addition to the COCO and NoCaps validation set, we evaluate the Out-domain performance of the model using Vizwiz caption dataset (Gurari et al., 2020). We use the same metrics as in Section 3.1.
@@
As shown in Table 3, incorporating sampled retrieved captions into training consistently enhances performance across various $k$ values. The improvement is particularly notable when captions are randomly retrieved, suggesting the model is stronger at ignoring irrelevant context. Improved cross-domain evaluation performance observed in Vizwiz is also consistent with the benefits of sampled training demonstrated in cross-domain machine translation by Hoang et al. (2022). Additionally, when comparing across different values of $k$, sampling mitigates the model’s sensitivity to the number of retrieved captions, outperforming top-$k$ training. For instance, it achieves comparable performance with a smaller $k$ value than in the case of top-$k$ training. Furthermore, the disparity between utilizing top-relevant and low-ranked retrieved captions is reduced with sample-$k$ training. Specifically, the maximum gap in CIDEr scores is reduced from 3 to 1, indicating increased model robustness even with less-relevant retrieved captions. Qualitative examples are displayed in Figure 6.
@@
Table 4: Training with sampled retrieved captions always outperforms SmallCap for all $k$ values for Vizwiz and NoCaps. The gain is smaller when using a larger datastore but still improves out-domain performance when retrieving more captions. Improved scores are in bold.
@@
We then evaluate on Vizwiz and Nocaps for its cross-domain performance (Table 4). This is a more realistic setting where retrieved captions are out-of-domain and could be more noisy and less relevant. The application of sampling improves across all $k$s for Vizwiz. With the COCO datastore, for Nocaps, sampling consistently improves near and out-domain performance, suggesting increased robustness to noisy retrieval context. Utilizing a broader data repository that incorporates internet-derived captions consistently improves in-domain performance. Retrieval constraints are alleviated for near and out-domain samples with the larger datastore, where we see smaller gains with sample-$k$. See qualitative examples in Figure 11 in Appendix C.
@@ 6 Discussion
To understand why the model depends heavily on majority tokens when making a prediction, we calculate how often majority tokens in the retrieved captions overlap with the ground truth captions ($T_{M}\in GT$) and with the predicted tokens ($T_{M}\in Pred$). From Table 5, we can observe that in 88%–99%of the training examples, the majority tokens in the retrieved captions are also present in the ground truth caption. This indicates that the model tends to depend on majority tokens, given their higher likelihood of being present in the ground truth label during training. This observation clarifies the decrease in the model’s robustness as $k$ increases when randomly retrieving captions. This is because a higher $k$ only adds noise without providing useful majority token. The introduction of sampling exposes the model to more diverse context during training, which leads to a slightly increased level of selectivity.
k 2 3 4
@@
Figure 7: Distribution of number of majority tokens in the retrieved captions for different evaluation datasets.
In Figure 7, we show the variation in the distribution of majority tokens across various evaluation datasets. When captions are randomly selected for the coco evaluation set, there are fewer majority token in the retrieved captions. This presents a challenge for the model in making use of the retrieved captions and generate accurate captions, which accounts for the performance decrease shown in Figure 1. For evaluation, with the same $k$, the fewer majority tokens exist in retrieved captions, the harder for the model to “cheat” by copy-and-past the majority tokens to the prediction. In such senarios, we obtain more significant improvements with the sample-$k$ training.
@@ 7 Conclusion and Future Work
We studied the robustness of the state-of-the-art retrieval-augmented image captioning model SmallCap and provide an through analysis and explanation of how retrieved captions effect the final prediction. Our exploration shows that SmallCap is robust to the order of the retrived captions, but it is sensitive to retrieval noise, which has implications for using retrieval-augmented models in new domains. With extensive input attribution analysis, we show that such sensitivity is due to majority tokens in the retrieved captions. We demonstrate a more retrieval robust model can be trained with sampling methods during training. We expect that our analysis can inspire better retrieval-robust captioning models in the future.
In the future, we will investigate whether the majority voting mechanism is exploited in other retrieval-augmented captioning models. We hope to further explore if other techniques such as token-dropping or prefix-tuning would further improve retrieval robustness.
@@
Despite the exploration in our work, we acknowledge that no system is perfect, and undesirable biases may still be present with our methods. We emphasize the need for continued research into techniques for identifying and mitigating hallucination and bias in retrieval-augmented models (Foulds et al., 2024; Deng et al., 2024). We also stress the importance of responsible deployment, with human oversight and content moderation pipelines in place.
@@ A.1 Stop words list
[’out’, ’some’, ’of’, ’is’, ’while’, ’are’, ’with’, ’down’, ’has’, ’over’, ’the’, ’next’, ’up’, ’near’, ’several’, ’other’, ’at’, ’top’, ’from’, ’in’, ’on’, ’a’, ’there’, ’an’, ’to’, ’and’, ’her’, ’front’, ’by’, ’for’, ’his’, ’it’]
@@ Appendix B More Visualization
Figure 8: Input attribution for each generated token (y-axis). Here the model prediction is misled by the majority tokens in the 2B1G setting.
@@ Appendix D More results
In Table 9 we show the percentage of tokens that are likely to be copied from retrieved captions averaging through all samples in the validation set. Majority tokens takes more than half of the copied tokens. With sample-$k$ training, model is slightly more selective when generating the captions.
k 1 2 3 4
@@
Table 10: CIDEr on COCO and NoCaps evaluation.

### v3

@@ 4.4 Attention and Model Behavior
Finally, we visualize the self-attention and cross-attention to locate the heads and layers in the SmallCap-OPT125M model that may contribute to the majority voting behaviour when generating a caption. This is crucial because all interactions between captions (self-attention) and images (cross-attention) take place in this stage.
@@
We partition the text input prompt into five distinct segments: begin of the sentence token (<BOS>), prompt tokens before retrieved $k$ captions (prefix), i.e. “Similar image shows”, the retrieved captions (retrieval) $cap_{1},\cdots cap_{k}$, prompt tokens before generation (suffix), i.e. “This image show”, and the generation itself. For image patches, we segment them into two pieces – the CLS output embedding, and the set of patch output embeddings.
Let $S_{n}$ denote the sets of indices, where $n=1,2,\ldots,5$ for five segments. For the text input, each segment $S_{n}$ contains the indices of the tokens in each segment. To track the occurrence of max attention values in $S_{n}$, we define the indicator function $\mathbbm{1}[I_{n}(i,j)]$ as follows:
$\mathbbm{1}[I_{n}(i,j)]=\begin{cases}1&\text{if }\arg\max_{z}Att(j,z)_{i}\in S_{n}\\ 0&\text{otherwise}\end{cases},$ (2)
where $\arg\max_{z}Att(j,z)_{i}$ is the index of the input with the maximum attention score for sample $i$.
For self-attention between the textual tokens, $Att(j,z)$ represents the attention score between the $j^{th}$ generated token and the $z^{th}$ text context token, denoted as $SA_{text}(j,z)$.
For cross-attention between the decoder and the image representations, we report both a text-centric and an image-centric analysis. The text-centric analysis $XA_{text}(j,z)$ measures the attention between the $j^{th}$ image patch and the $z^{th}$ text token, to identify which segments of the text have the highest cross-attention scores in relation to the image. In the image-centric analysis $XA_{img}(j,z)$, we measure the attention between the $j^{th}$ generated token and the $z^{th}$ image patch. We now redefine the $S_{n}$ notation to let $S_{1}$ represent the CLS output embedding, and $S_{2}$ represent the set of image patch embeddings, respectively. This allows $XA_{img}(j,z)$ to identify if the CLS patch embedding receives the highest cross-attention scores in relation to the generated tokens, or if it is the actual image patch embeddings.
For each analysis, $SA_{text}(j,z)$, $XA_{text}(j,z)$, and $XA_{img}(j,z)$, we calculate the proportion of occurrences of the maximum score in $S_{n}$ by averaging through all generated tokens for a dataset.
@@
We gather attention scores between the generated tokens and context tokens, and categorize the distribution of the maximum scores into the five text segments (BOS, prefix, retrieval, suffix, and generation).
Figure 5(a) illustrates the changes in the distribution of maximum self-attention scores in each layer of the decoder language model. Notably, at the initial layers, a majority of attention heads exhibit heightened focus on retrieved captions or the current context for generation. However, after the second layer, we observe an increased emphasis on the beginning of sentence token (</s>). This behavior is consistent with prior research on the attention mechanism of GPT-2 Vig and Belinkov (2019). Figures LABEL:subfig:gpt2-self-attention and LABEL:subfig:opt-self-attention show the behaviour for all self-attention heads in for the GPT and OPT model variants, respectively.
@@
Similar to the self-attention behaviour, we categorize the occurrence of the maximum cross-attention to the five text segments. As shown in Figure 5(b), in most attention heads, the cross-attention attains its maximum value between the image and the retrieved captions or between the image and the generated tokens. Figures LABEL:subfig:gpt2-cross-attention-text and LABEL:subfig:opt-cross-attention-text show the text-centric analysis for all cross-attention heads for the GPT and OPT backbones.
Finally, we inspect whether the model focuses on the CLS patch or actual image patches. In Figure 5(c), we observe that the model only pays maximum attention to the image patches in the final layers (the blue line). Figures LABEL:subfig:gpt2-cross-attention-cls and LABEL:subfig:opt-cross-attention-cls show the full results for the image-centric analysis.
Overall, these observations show that the model attends to both modalities during the caption generation process. However the lack of strong cross-attention to actual image patches suggests that the model is misled by text prompts, even when irrelevant information is absent in the provided image.
@@ 5 Improving Robustness to Retrieval via Sampling
In order to improve the robustness of the model to potentially noisy captions, we propose to randomly sample the captions from a larger retrieval list for a given image, instead of training with only the top-$k$ retrieved captions. In this manner, the model can learn from more diverse context that includes both top- and lower-ranked captions.
@@ 5.1 Experimental Setup
We sample $k$ captions randomly from the top-$N$=7 retrieved captions during training [^5] We sample from the top-$N$=7 for alignment with the baseline; see the Appendix for an ablation on varying $N$.. Following Ramos et al. (2023b), we train smallCap with the OPT-350M decoder on the COCO captioning dataset (Chen et al., 2015) for 10 epochs on a NVIDIA A100 40GB GPU with the default learning rate of 1e-4 and batch size of 64. We experiment with $k$ in the range of 1–4.
@@ 5.2 Evaluation and Results
In addition to the COCO and NoCaps validation set, we evaluate the Out-domain performance of the model using VizWiz caption dataset (Gurari et al., 2020) and report CIDEr scores.
COCO Eval
Model $k$ top-$k$ last-$k$ random
top-k 1 115.1 112.2 73.2
sample-k 1 116.0 115.0 98.9
top-k 2 116.8 115.0 67.4
sample-k 2 117.4 116.8 84.6
top-k 3 118.3 117.1 71.8
sample-k 3 118.5 117.3 77.6
top-k 4 120.1 117.1 70.1
sample-k 4 119.2 118.6 73.1
c-sample-k 4 119.3 118.9 72.6
Table 3: CIDEr scores when training on the top-$k$, sample-$k$ and c-sample-$k$ captions. Training by sampling the retrieved captions almost always outperforms SmallCap for all $k$ values. It also reduces the gap between using top-relevant and low-ranked retrieved captions. Results are averaged over three seeds. Improved scores are in bold.
@@
As shown in Table 3, incorporating sampled retrieved captions into training consistently enhances performance across various $k$ values. The improvement is particularly notable when captions are randomly retrieved, suggesting the model is now better able to ignore irrelevant context. If we compare across different values of $k$, sampling mitigates the model’s sensitivity to the number of retrieved captions, outperforming top-$k$ training. For instance, it achieves comparable performance with a smaller $k$ value than in the case of top-$k$ training. Furthermore, the gap between using the top-$k$ vs. the last-$k$ retrieved captions is reduced with sample-$k$ training: the maximum gap is reduced from 3.0 to 1.0 CIDEr points, indicating increased model robustness, even with lower-ranked retrieved captions. Figure 6 and 12 show qualitative examples of the improved robustness to randomly retrieved examples.
Figure 6: Qualitative examples of generated captions when randomly retrieving four captions for a given image using a model trained with either the Sample-k or the Top-k method.
@@
Table 4: Training with sampled retrieval always outperforms top-$k$ retrieval for all values of $k$ on the out-of-domain VizWiz and NoCaps datasets. The gains are smaller when using a larger datastore (+Web) but it still improves out-domain performance when retrieving more captions. Improved scores are in bold.
@@
We also evaluate on VizWiz and NoCaps to measure cross-domain performance (Table 4). This is a more realistic setting where retrieved captions are out-of-domain and could be more noisy and less relevant. The application of sampling improves across all values of $k$ for Vizwiz. On the NoCaps dataset, with the COCO datastore, sampling consistently improves near and out-domain performance, suggesting increased robustness to noisy retrieval context. This is consistent with the benefits of sampled training demonstrated in cross-domain machine translation by Hoang et al. (2022). If we use a larger datastore that incorporates internet-derived captions (+Web), this consistently improves in-domain performance. Retrieval constraints are alleviated for near and out-domain samples with the larger datastore, where we see smaller gains with sample-$k$. See qualitative examples in Figure 11 in Appendix C.
@@ 6 Discussion
To better understand why the model relies on majority tokens during generation, we calculate the probability that majority tokens in the retrieved captions overlap with the ground truth captions ($T_{M}\in GT$), and with the predicted tokens ($T_{M}\in Pred$). Table 5 shows that in 88%–99% of the training examples, the majority tokens in the retrieved captions are also present in the ground truth captions. This suggests that the model can develop a bias towards majority tokens due to the fact that they are so often present in the ground truth during training. This analysis also clarifies the decrease in the model’s robustness as $k$ increases when randomly retrieving captions. This is because a higher $k$ only adds noise without providing useful majority tokens. The use of sampling during training exposes the model to more diverse context, which leads to a slightly increased level of selectivity.
k=2 3 4
@@
Figure 7: Distribution of number of majority tokens in the retrieved captions for the COCO, VizWis, and NoCaps evaluation datasets. For the COCO dataset, we also show the difference between retrieving the top-4 captions against four randomly selected captions.
In Figure 7, we show the variation in the distribution of majority tokens across various evaluation datasets. When captions are randomly selected for the COCO evaluation data, there are fewer majority tokens in the retrieved captions. This presents a challenge for the model in making use of the retrieved captions, which accounts for the performance decrease shown in Figure 1. For evaluation, with the same value of $k$, the fewer the number of majority tokens in the retrieved captions, the harder it is for the model to “copy” those tokens to the final output. In such scenarios, we obtain bigger improvements with the sample-$k$ training.
@@ 7 Conclusion and Future Work
We studied the robustness of the state-of-the-art retrieval-augmented image captioning model SmallCap and provide an through analysis and explanation of how retrieved captions effect the final prediction. Our exploration shows that SmallCap is robust to the order of the retrieved captions, but it is sensitive to retrieval noise, which has implications for using retrieval-augmented models in new domains. With extensive input attribution analysis, we show that such sensitivity is due to majority tokens in the retrieved captions. We demonstrate a more retrieval robust model can be trained with sampling methods during training. We expect that our analysis can inspire better retrieval-robust captioning models in the field.
In the future, we will investigate whether the majority voting behaviour is exploited in other retrieval-augmented captioning models. We hope to further explore if other techniques such as token-dropping or prefix-tuning would further improve retrieval robustness.
@@
Despite the exploration in our work, we acknowledge that no system is perfect, and undesirable biases may still be present with our methods. We emphasize the need for continued research into techniques for identifying and mitigating hallucination and bias in retrieval-augmented models (Foulds et al., 2024; Deng et al., 2024). We also stress the importance of responsible deployment, with human oversight and content moderation pipelines.
@@ A.1 Stop words list
[“out”, “some”, “of”, “is”, “while”, “are”, “with”, “down”, “has”, “over”, “the”, “next”, “up”, “near”, “several”, “other”, “at”, “top”, “from”, “in”, “on”, “a”, “there”, “an”, “to”, “and”, “her”, “front”, “by”, “for”, “his”, “it”]
@@ Appendix B More Visualization
Figure 8: Attribution visualization with few more examples. Here the model prediction is misled by the majority tokens in the 2B1G setting.
@@ Appendix D More results
In Table 9 we show the percentage of tokens that are likely to be copied from retrieved captions averaging through all samples in the validation set. Majority tokens takes more than half of the copied tokens.
k=1 2 3 4
@@
Table 10: CIDEr on COCO and NoCaps.